# Proactive Model Adaptation Against Concept Drift for Online Time Series Forecasting

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 3–7, 2025; Toronto, Canada.isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data stream miningccs: Computing methodologies Online learning settings
@@ 1. Introduction
Time series forecasting has been a prevalent task in numerous fields such as energy (Boussif et al., 2023), retail (Böse et al., 2017), climate (Bi et al., 2023) and finance (Zhao et al., 2023). Recent years have witnessed a surge of deep learning-based forecast models (Nie et al., 2023; Liu et al., 2024; Zhou et al., 2023; Woo et al., 2024; Han et al., 2024) that take past time series observations to predict values in the next $H$ steps, where $H$ is called forecast horizon. Due to the dynamic nature of the environment, latent concepts that influence observation values (e.g., social interest (Bai et al., 2023), stock market sentiment (Li et al., 2022)) often change over time. This ubiquitous phenomenon is known as concept drift (Gama et al., 2014). In the presence of concept drift, future test data may not follow a similar data distribution as the historical training data, causing degradation in forecasting performance (You et al., 2021; Zhao et al., 2023).
@@
As illustrated in Fig. 1, an online learning task inherently has a $H$-step feedback delay in time series forecasting, resulting in a temporal gap (at least $H$ steps) between available training samples and the test sample. Formally, at each online time $t$, the current horizon window $\mathbf{Y}_{t}$, which have not been observed yet, has overlap with $\mathbf{Y}_{t-H+1},\dots,\mathbf{Y}_{t-1}$. These samples cannot provide complete forecasting feedback and supervisions for the performant, prevailing time series forecasting models which predict $H$-step values directly (Zeng et al., 2023; Nie et al., 2023; Liu et al., 2024). Hence, the available training samples for online learning at time $t$ is $\mathcal{D}_{t-}=\{(\mathbf{X}_{t^{\prime}},\mathbf{Y}_{t^{\prime}})\mid t^{\prime}\leq t-H\}$, which has a temporal distance to the test sample $(\mathbf{X}_{t},\mathbf{Y}_{t})$. The temporal gap issue suggests that addressing concept drift by fitting the forecast model to the recent training samples may be insufficient, since concept drift may also occur over the $H$-step temporal gap between the recent training samples and the current test sample.
@@
To investigate the impact of this temporal gap on forecasting performance, we examine two online learning strategies using real-world time series datasets (detailed in Section 3). The first strategy fine-tunes model parameters using the latest available training sample at time $t$, which is $(\mathbf{X}_{t-H},\mathbf{Y}_{t-H})$. The second strategy (Pham et al., 2023; Zhang et al., 2023) omits the feedback delay and adapts the model using $(\mathbf{X}_{t-1},\mathbf{Y}_{t-1})$, which is infeasible in practice due to future information leakage. The empirical results show that the average forecast error of the first strategy is approximately double that of the second strategy, indicating that even the latest observed time series pattern differs from the test sample notably. Moreover, the performance gap between the two strategies becomes more significant with a longer horizon, suggesting that the concept drift becomes more pronounced over longer time intervals. These findings highlight the presence of a substantial concept drift between the practical training samples and the test sample, limiting the effectiveness of current online learning techniques that adapt model parameters to the potentially outdated concepts and leave the concept drift unresolved.
In this paper, we aim to answer the question: how can we effectively adapt the forecast model to each test sample and boost online time series forecasting performance against ongoing concept drifts? A straightforward idea is to estimate a sketch of the latent concept of the test data and customize model parameters for the estimated concept. However, it is challenging to train a model adapter (e.g., a mapping function) that directly generates model parameters based on the estimated concept of the test sample. First, test data may bear a new concept that is out of the distribution of the historical time series. There is no experience in generating optimal parameters for such an out-of-distribution concept. Second, a parameterized model adapter that maps concepts to model parameters (or their updates) essentially lies in a parameter space of $\mathbb{R}^{m\times n}$, where $m$ is the concept embedding dimension and $n$ is the number of model parameters. Since advanced time series forecasting models often involve a high quantity of parameters (e.g., 1 million (Nie et al., 2023; Liu et al., 2024)), the model adapter is hard to optimize and easily suffer from overfitting due to the huge parameter size and limited test data.
@@ 3.3. Evaluation Metrics
Table 1. MSE results of two online learning strategies for time series forecasting. The horizon $H$ varies in 24, 48, and 96.
@@ 4.1. Solution Overview
    Online forecasting. Finally, the adapted model yields predictions by $\hat{\mathbf{Y}}_{t}=\mathcal{F}(\mathbf{X}_{t};\hat{\boldsymbol{\theta}}_{t})$. In the next time $t+1$, the parameters will be reset to $\boldsymbol{\theta}_{t-H}$.
@@
The rationale of our solution is that we can synthesize diverse concept drifts based on historical data to train our model adapter. Fig. 3 shows an example of the historical training data including four samples with their concepts denoted by $\mathbf{c}_{1},\cdots,\mathbf{c}_{4}$ respectively. Our idea is to shuffle the order of samples and train our model adapter on the synthetic concept drifts between each pair of samples. Though the concepts of the future test samples (i.e., $\mathbf{c}_{5},\cdots,\mathbf{c}_{7}$) are out-of-distribution, the concept drift patterns $\boldsymbol{\delta}_{4\to 6}$ and $\boldsymbol{\delta}_{5\to 7}$ are similar to $\boldsymbol{\delta}_{1\to 4},\boldsymbol{\delta}_{2\to 3}$ which have been learned in training data. Given such reoccuring concept drifts, we can generate appropriate parameter shifts by experience.
@@ 4.2. Concept Drift Estimation
It is noteworthy that $\mathcal{E}^{\prime}$ has the potential to estimate hidden state of $\mathbf{Y}_{t}$. When $L\geq kH$ and $k\geq 2$, $\mathbf{X}_{t}$ itself contains a sequence of horizon windows, i.e., $\{\mathbf{Y}_{t-kH},\cdots,\mathbf{Y}_{t-H}\}$. In this case, $\mathtt{MLP^{\prime}}$ can learn the hidden state of each observed horizon windows, exploit the temporal evolution pattern across horizon windows, and extrapolate the hidden state of the next horizon window $\mathbf{Y}_{t}$.
@@ 4.5. Discussion
Prior approaches to online time series forecasting make model adaptation passively based on feedbacks in forecasting previous samples, which mainly focus on learning recent data patterns. In Table 2, we compare existing methods and Proceed in terms of the training data and the model adaptation techniques at each online time $t$. Most methods use one newly acquired training sample, while SOLID (Chen et al., 2024) selects more recent samples that share similar lookback windows with the test sample and are assumed to share similar concept. SOLID and Proceed simply fine-tune the forecast model by gradient descent, while FSNet and OneNet further generates additional parameter adjustments based on the forecasting feedback. Since the feedback is delayed by $H$ steps and cannot reveal the test concept, we propose a novel step called proactive model adaptation which aims to mitigate the effects of concept drift between the training sample and the test sample. As this core step is orthogonal to existing methods, Proceed can incorporate the data augmentation technique of SOLID and the feedback-based model adaptation techniques of FSNet and OneNet.
@@ 5. Experiments
5.1. Experimental Settings
[…]
Datasets. As introduced in Sec. 3.1, we use five popular benchmarks, of which more details are provided in Appendix C.1. We conform to the dataset split ratio of FSNet and OneNet, i.e., we split the datasets by 20:5:75 for training, validation, and testing. The rationale is that online learning is of great practical value in scenarios of limited training data, when pretrained forecast models are inadept at handling new concepts during long-term online service.
[…]
Online Learning Baselines. We compare Proceed with GD, a naïve online gradient descent method, and the state-of-the-art online model adaptation methods, including FSNet, OneNet, and SOLID. We reimplement FSNet and OneNet by adopting the Practical strategy described in Sec. 3.2. Furthermore, FSNet is specially designed for TCN, while OneNet is based on online ensembling and is model-agnostic. We implement multiple variants of OneNet by combining each forecast model with an FSNet. As for SOLID, its original work learns linear probing with the pretrained parameters at each update time and does not inherit the fine-tuned parameters from the last update. As our datasets have a much long online phase, we implement a variant called SOLID++ that continually fine-tunes all model parameters across the online data. Empirically, SOLID++ performs better than SOLID in our evaluation setting.
@@ 5.2.2. Efficiency
As depicted in Fig. 5, we compare the GPU memory occupation and inference latency of different model adaptation methods on Traffic, the largest time series dataset. We record the online latency, including feedback-based model adaptation and any other additional steps. Here, SOLID only fine-tunes the final layer rather than the whole model, reducing the gradient computation cost. More specifically, SOLID and SOLID++ computes gradients on several selected training samples iteratively instead of concurrently, so as to save GPU memory. By contrast, Proceed only uses the latest training sample and thereby achieves the lowest latency. Moreover, Proceed is faster and more lightweight than the online ensembling method OneNet.
@@ 5.3. Visualization of the Representation Space
To verify our assumption that online data have OOD concepts, we adopt t-SNE to visualize the representations of concepts $\mathbf{c}_{t}$ and concept drifts $\boldsymbol{\delta}_{t-H\to t}$ on the historical training data and online data. As shown in Fig. 6(a), there are a great number of OOD concepts on the online data, which are distinct from any historical concept. By contrast, as shown in Fig. 6(b), we encounter much fewer OOD concept drifts on the online data. These results support our intuition illustrated in Fig. 3. Hence, it is desirable for the adaptation generator to generate adaptations based on concept drifts instead of concepts.
@@ 6. Conclusion
In this work, we highlight that online time series forecasting inherently has a temporal gap between each test sample and available training data, where concept drift may well occur. Through empirical study, we found that this gap hinders the effectiveness of existing model adaptation methods which passively rely on the feedback in forecasting recent data. To address this issue, we propose a novel online model adaptation framework named Proceed, which proactively adapts the forecast model to the test concept before forecasting each test sample. Specifically, Proceed first fine-tunes the model on the latest acquired training sample, then extracts latent features from time series data to estimate the undergoing concept drift, and efficiently maps the estimated drift into parameter adjustments that are tailored for the test sample. Furthermore, we synthesize diverse concept drift and optimize Proceed with generalization capacity of mapping concept drift to beneficial parameter adjustments. Extensive experiments on five real-world time series datasets demonstrate that our proposed Proceed remarkably reduces the forecast errors of various forecast models and surpasses the state-of-the-art online learning methods. Our proactive model adaptation method provides a new direction in addressing continuous concept drift for online systems, which we wish to help any other prediction tasks with a feedback delay issue.
@@ B.1. Online Model Adaptation
After the submission of our manuscript, there are two concurrent researches (Anonymous, 2024; Liang et al., 2024) that also noticed the information leakage in FSNet and OneNet. Since the ground-truth $H$-step values of the last prediction are not fully observed, both works propose to generate pseudo labels of the unobserved values, concatenate the pseudo labels with observed labels to simulate complete ground truth, and calculate the forecast errors to update the model. The potential drawback is that the pseudo label generator is still susceptible to concept drift and may produce low-quality pseudo labels, leaving the concept drift issue unresolved.
@@ Appendix D Pipeline of Proceed
Fig. 7 depicts the pipeline of Proceed during the online phase, which is a bit different from the steps introduced in Sec. 4.1. Given model parameters $\boldsymbol{\theta}_{t-H-1}$ that have been updated with previous training data $\mathcal{D}_{(t-1)-}$, our model adapter additionally estimates the concept drift between $\mathcal{D}_{(t-1)-}$ and $\mathcal{D}_{t-}$ and adapts the model to $\mathcal{D}_{t-}$, yielding forecast feedback for updates. The reason behind it is that we jointly train the forecast model and the model adapter (see Sec. 4.4), while our training algorithm does not enforce the forecast model to perform well alone. Consequently, the forecast model without $\Delta\boldsymbol{\theta}$ could forget some predictive skills, while the parameter adjustments $\Delta\boldsymbol{\theta}$ from the model adapter may incorporate the skills forgot by the model. Thus it is necessary for the inference process to keep consistent with the training phase, i.e., we always make predictions by integrating the forecast model and our model adapter as a whole. Nevertheless, in the feedback-based adaptation step, we only finetune the forecast model but keep the model adapter’s parameters frozen. Otherwise, the model adapter may overfit the patterns of the one-step concept drift between $\mathcal{D}_{(t-1)-}$ and $\mathcal{D}_{t-}$, losing generalization ability.

### v3

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1245-6/25/08doi: 10.1145/3690624.3709210 ccs: Information systems Data stream miningccs: Computing methodologies Online learning settings
@@ 1. Introduction
Time series forecasting has been a prevalent task in numerous fields such as climate (Bi et al., 2023), energy (Boussif et al., 2023; Sun et al., 2023), retail (Böse et al., 2017), and finance (Zhao et al., 2023). Recent years have witnessed a surge of deep learning-based forecast models (Nie et al., 2023; Liu et al., 2024; Zhou et al., 2023; Woo et al., 2024; Han et al., 2024) that take past time series observations to predict values in the next $H$ steps, where $H$ is called forecast horizon. Due to the dynamic nature of the environment, latent concepts that influence observation values (e.g., social interest (Bai et al., 2023), stock market sentiment (Li et al., 2022)) often change over time. This ubiquitous phenomenon is known as concept drift (Gama et al., 2014). In the presence of concept drift, future test data may not follow a similar data distribution as the historical training data, causing degradation in forecasting performance (You et al., 2021; Zhao et al., 2023).
@@
As illustrated in Fig. 1, an online learning task inherently has a $H$-step feedback delay in time series forecasting, resulting in a temporal gap (at least $H$ steps) between available training samples and the test sample. Formally, at each online time $t$, the current horizon window $\mathbf{Y}_{t}$, which has not been observed yet, has an overlap with $\mathbf{Y}_{t-H+1},\dots,\mathbf{Y}_{t-1}$. These samples cannot provide complete forecasting feedback and supervision for the performant, prevailing time series forecasting models that predict $H$-step values directly (Zeng et al., 2023; Nie et al., 2023; Liu et al., 2024). Hence, the available training samples for online learning at time $t$ is $\mathcal{D}_{t-}=\{(\mathbf{X}_{t^{\prime}},\mathbf{Y}_{t^{\prime}})\mid t^{\prime}\leq t-H\}$, which has a temporal distance to the test sample $(\mathbf{X}_{t},\mathbf{Y}_{t})$. The temporal gap issue suggests that addressing concept drift by fitting the forecast model to the recent training samples may be insufficient, since concept drift may also occur over the $H$-step temporal gap between the recent training samples and the current test sample.
@@
To investigate the impact of this temporal gap on forecasting performance, we examine two online learning strategies using real-world time series datasets (detailed in Section 3). The first strategy fine-tunes model parameters using the latest available training sample at time $t$, which is $(\mathbf{X}_{t-H},\mathbf{Y}_{t-H})$. The second strategy (Pham et al., 2023; Zhang et al., 2023) omits the feedback delay and adapts the model using $(\mathbf{X}_{t-1},\mathbf{Y}_{t-1})$, which is infeasible in practice due to future information leakage. The empirical results show that the average forecast error of the first strategy is approximately double that of the second strategy, indicating that even the latest observed time series pattern differs from the test sample notably. Moreover, the performance gap between the two strategies becomes more significant with a longer horizon, suggesting that the concept drift becomes more pronounced over longer time intervals. These findings highlight the presence of a substantial concept drift between the practical training samples and the test sample, limiting the effectiveness of current online learning techniques that adapt model parameters to potentially outdated concepts and leave the concept drift unresolved.
In this paper, we aim to answer the question: how can we effectively adapt the forecast model to each test sample and boost online time series forecasting performance against ongoing concept drifts? A straightforward idea is to estimate a sketch of the latent concept of the test data and customize model parameters for the estimated concept. However, it is challenging to train a model adapter (e.g., a mapping function) that directly generates model parameters based on the estimated concept of the test sample. First, test data may bear a new concept that is out of the distribution of the historical time series. There is no experience in generating optimal parameters for such an out-of-distribution concept. Second, a parameterized model adapter that maps concepts to model parameters (or their updates) essentially lies in a parameter space of $\mathbb{R}^{m\times n}$, where $m$ is the concept embedding dimension and $n$ is the number of model parameters. Since advanced time series forecasting models often involve a high quantity of parameters (e.g., 1 million (Nie et al., 2023; Liu et al., 2024)), the model adapter is hard to optimize and easily suffers from overfitting due to the huge parameter size and limited test data.
@@ 3.3. Evaluation Metrics
Table 1. MSE results of two online learning strategies for time series forecasting. We report the MAE results in Appendix E.1.
@@ 4.1. Solution Overview
    Online forecasting. Finally, the adapted model yields predictions by $\hat{\mathbf{Y}}_{t}=\mathcal{F}(\mathbf{X}_{t};\hat{\boldsymbol{\theta}}_{t})$. At the next time $t+1$, the parameters will be reset to $\boldsymbol{\theta}_{t-H}$.
@@
The rationale of our solution is that we can synthesize diverse concept drifts based on historical data to train our model adapter. Fig. 3 shows an example of the historical training data including four samples with their concepts denoted by $\mathbf{c}_{1},\cdots,\mathbf{c}_{4}$ respectively. Our idea is to shuffle the order of samples and train our model adapter on the synthetic concept drifts between each pair of samples. Though the concepts of the future test samples (i.e., $\mathbf{c}_{5},\cdots,\mathbf{c}_{7}$) are out-of-distribution, the concept drift patterns $\boldsymbol{\delta}_{4\to 6}$ and $\boldsymbol{\delta}_{5\to 7}$ are similar to $\boldsymbol{\delta}_{1\to 4},\boldsymbol{\delta}_{2\to 3}$ which have been learned in training data. Given such recurring concept drifts, we can generate appropriate parameter shifts by experience.
@@ 4.2. Concept Drift Estimation
It is noteworthy that $\mathcal{E}^{\prime}$ has the potential to estimate the hidden state of $\mathbf{Y}_{t}$. When $L\geq kH$ and $k\geq 2$, $\mathbf{X}_{t}$ itself contains a sequence of horizon windows, i.e., $\{\mathbf{Y}_{t-kH},\cdots,\mathbf{Y}_{t-H}\}$. In this case, $\mathtt{MLP^{\prime}}$ can learn the hidden state of each observed horizon window, exploit the temporal evolution pattern across horizon windows, and extrapolate the hidden state of the next horizon window $\mathbf{Y}_{t}$.
@@ 4.5. Discussion
Prior approaches to online time series forecasting perform model adaptation based on feedback in forecasting previous samples, which mainly focus on learning recent data patterns. In Table 2, we compare existing methods and Proceed in terms of the training data and the model adaptation techniques at each online time $t$. Most methods use one newly acquired training sample, while SOLID (Chen et al., 2024) selects more recent samples that share similar lookback windows with the test sample and are assumed to share a similar concept. SOLID and Proceed simply fine-tune the forecast model by gradient descent, while FSNet and OneNet further generate additional parameter adjustments based on the forecasting feedback. Since the feedback is delayed by $H$ steps and cannot reveal the test concept, we propose a novel step called proactive model adaptation which aims to mitigate the effects of concept drift between the training sample and the test sample. As this core step is orthogonal to existing methods, Proceed can incorporate the data augmentation technique of SOLID and the feedback-based model adaptation techniques of FSNet and OneNet.
@@ 5. Experiments
5.1. Experimental Settings
[…]
Datasets. As introduced in Sec. 3.1, we use five popular benchmarks, of which more details are provided in Appendix C.1. We conform to the dataset split ratio of FSNet and OneNet, i.e., we split the datasets by 20:5:75 for training, validation, and testing. The rationale is that online learning is of great practical value in scenarios of limited training data, when pretrained models are inadequate at handling new concepts during long-term online service.
[…]
Online Learning Baselines. We compare Proceed with GD, a naïve online gradient descent method, and the state-of-the-art online model adaptation methods, including FSNet, OneNet, and SOLID. We reimplement FSNet and OneNet by adopting the Practical strategy described in Sec. 3.2. Furthermore, FSNet is specially designed for TCN, while OneNet is based on online ensembling and is model-agnostic. We implement multiple variants of OneNet by combining each forecast model with an FSNet. As for SOLID, its original work learns linear probing with the pretrained parameters at each update time and does not inherit the fine-tuned parameters from the last update. As our datasets have a much longer online phase, we implement a variant called SOLID++ that continually fine-tunes all model parameters across the online data. Empirically, SOLID++ performs better than SOLID in our evaluation setting.
@@ 5.2.2. Efficiency
As depicted in Fig. 5, we compare the GPU memory occupation and inference latency of different model adaptation methods on Traffic, the largest time series dataset. We record the online latency, including feedback-based model adaptation and any other additional steps. Here, SOLID only fine-tunes the final layer rather than the whole model, reducing the gradient computation cost. More specifically, SOLID and SOLID++ compute gradients on several selected training samples iteratively instead of concurrently, so as to save GPU memory. By contrast, Proceed only uses the latest training sample and thereby achieves the lowest latency. Moreover, Proceed is faster and more lightweight than the online ensembling method OneNet.
@@ 5.3. Visualization of the Representation Space
To verify our assumption that online data have OOD concepts, we adopt t-SNE to visualize the representations of concepts $\mathbf{c}_{t}$ and concept drifts $\boldsymbol{\delta}_{t-H\to t}$ on the historical training data and online data. As shown in Fig. 6(a), there are a great number of OOD concepts in the online data, which are distinct from any historical concept. By contrast, as shown in Fig. 6(b), we encounter much fewer OOD concept drifts on the online data. These results support our intuition illustrated in Fig. 3. Hence, it is desirable for the adaptation generator to generate adaptations based on concept drifts instead of concepts.
@@ 5.4. Ablation Study
Besides, we study different kinds of operations on multivariate time series using our concept encoder. Let $\mathbf{c}_{t}^{(i)}$ denote the latent concept feature vector of the $i$-th variate. Apart from the average introduced in Eq. (1), we implement two variants as follows:
  • linear transformation: we learn a linear layer that transforms the concatenation of all latent features, i.e., $c_{t}=W^{\top}[c_{t}^{(0)},c_{t}^{(1)},\cdots]$.
  • weighted sum: we learn an individual weight $w_{i}\in\mathbb{R}$ for each $i$-th variate, and $c_{t}=\sum_{i}w_{i}c_{t}^{(i)}$.
In Table 6, we compare the performance of the variants in terms of the average MSE with $H\in\{24,48,96\}$. The results demonstrate that the average operation is simple yet effective. One possible reason is that all variates are typically assumed to have equal importance during both training and evaluation.
Table 6. MSE of Proceed with various concept encoders.
Model Method ETTh2 ETTm1 Weather ECL Traffic
PatchTST linear trans. 3.498 0.570 0.985 5.661 0.363
weighted sum 3.467 0.571 0.986 5.435 0.362
average 3.468 0.554 0.986 4.732 0.356
iTransfo. linear trans. 4.400 0.548 1.018 4.944 0.355
weighted sum 4.441 0.551 1.021 5.260 0.352
average 4.216 0.543 1.017 4.752 0.352
@@ 6. Conclusion
In this work, we highlight that online time series forecasting inherently has a temporal gap between each test sample and available training data, where concept drift may well occur. Through empirical study, we found that this gap hinders the effectiveness of existing model adaptation methods which passively rely on the feedback in forecasting recent data. To address this issue, we propose a novel online model adaptation framework named Proceed, which proactively adapts the forecast model to the test concept before forecasting each test sample. Specifically, Proceed first fine-tunes the model on the latest acquired training sample, then extracts latent features from time series data to estimate the undergoing concept drift, and efficiently maps the estimated drift into parameter adjustments that are tailored for the test sample. Furthermore, we synthesize diverse concept drift and optimize Proceed with the generalization capacity of mapping concept drift to beneficial parameter adjustments. Extensive experiments on five real-world time series datasets demonstrate that our proposed Proceed remarkably reduces the forecast errors of various forecast models and surpasses the state-of-the-art online learning methods. Our proactive model adaptation method provides a new direction in addressing continuous concept drift for online systems, which we wish to help any other prediction tasks with a feedback delay issue.
@@ B.1. Online Model Adaptation
After the submission of our manuscript, there are two concurrent researches (yee Ava Lau et al., 2025; Liang et al., 2024) that also noticed the information leakage in FSNet and OneNet. Since the ground-truth $H$-step values of the last prediction are not fully observed, both works propose to generate pseudo labels of the unobserved values, concatenate the pseudo labels with observed labels to simulate complete ground truth, and calculate the forecast errors to update the model. The potential drawback is that the pseudo label generator is still susceptible to concept drift and may produce low-quality pseudo labels, leaving the concept drift issue unresolved.
@@ Appendix D Pipeline of Proceed
Fig. 7 depicts the pipeline of Proceed during the online phase, which is a bit different from the steps introduced in Sec. 4.1. Given model parameters $\boldsymbol{\theta}_{t-H-1}$ that have been updated with previous training data $\mathcal{D}_{(t-1)-}$, our model adapter additionally estimates the concept drift between $\mathcal{D}_{(t-1)-}$ and $\mathcal{D}_{t-}$ and adapts the model to $\mathcal{D}_{t-}$, yielding forecast feedback for updates. The reason behind it is that we jointly train the forecast model and the model adapter (see Sec. 4.4), while our training algorithm does not enforce the forecast model to perform well alone. Consequently, the forecast model without $\Delta\boldsymbol{\theta}$ could forget some predictive skills, while the parameter adjustments $\Delta\boldsymbol{\theta}$ from the model adapter may incorporate the skills forgotten by the model. Thus it is necessary for the inference process to keep consistent with the training phase, i.e., we always make predictions by integrating the forecast model and our model adapter as a whole. Nevertheless, in the feedback-based adaptation step, we only finetune the forecast model but keep the model adapter’s parameters frozen. Otherwise, the model adapter may overfit the patterns of the one-step concept drift between $\mathcal{D}_{(t-1)-}$ and $\mathcal{D}_{t-}$, losing generalization ability.